42. Three substances have the following specific heat capacities: Substance A: 1 cal/goC
Substance B: 0.50 cal/goC

Substance C: 0.25 cal/goC
Which substance will experience the largest temperature change if 100 kJ of energy are applied to each one of them and we have the same mass of each substance?
Answer #1

Q = m c ∆T
Q = quantity of heat in joules (J)
m = mass of the substance acting as the environment in
grams (g)
c = specific heat capacity in J/(g oC)
∆T = change in temperature = Tfinal - Tinitial in oC

∆T = Q / m c

Substance A: 1 cal/goC or 4.184 J/(g oC)

If we keep Mass of the substance = 1000 g

∆T = 100 x 103 / 1000 x 4.184 = 23.9 oC

Substance b: 0.5 cal/goC or 2.092 J/(g oC)

∆T = 100 x 103 / 1000 x 2.092 = 47.8   oC

Substance b: 0.25 cal/goC or 1.046 J/(g oC)

∆T = 100 x 103 / 1000 x 1.046 = 95.6 oC

Hence substance C will experience the largest temperature change (95.6 oC)